OpenStudy (anonymous):

Maths Problem: Three nuts and six bolts have a combined mass of 72g. Four nuts and 5 bolts have a combined mass of 66g. Find the combined mass of one nut and one bolt

OpenStudy (anonymous):

3n+6b=72 4n+5b=66 now multiply equation one by 4 and equation 2 by 3 12n+24b=288 12n+15b=198 now subtract equation one by 2 9b=90 b=10 put b=10 in any equation 3n+6*10=72 3n=72-60 3n=12 n=4 therefore nuts=4 and bolts =10

OpenStudy (anonymous):

hope that helps ya miss chloris, by the way thers an option to give medals lol

OpenStudy (anonymous):

LOL :D

OpenStudy (anonymous):

keep rocking ;)

OpenStudy (anonymous):

How about these? :(

OpenStudy (anonymous):

its too small i cant see anything

OpenStudy (anonymous):

How about now?

OpenStudy (anonymous):

awesome , ya rock, hold on a sec

OpenStudy (anonymous):

length and width are equal to each other therefore 4a=3b+1 2a-3=5-b 4a-3b=1 2a+b=8 multiply second eqation by 2 4a-3b=1 4a+2b=16 now subtract equation 1 from 2 -5b=-15 b=3 and a= 5/2